# Whirling injections, surjections, and other functions between finite sets

### v5

@@ Abstract.
Keywords: dynamical algebraic combinatorics, homomesy, injection, noncrossing partition, RG-word, set partition, surjection, toggles, toggle group, whirling
1. Introduction
In this paper, we explore the orbits associated with an action that can be defined on any family $\mathcal{F}$ of functions $f$ from one finite set $S$ to another finite set $T$, where $S$ comes with a linear ordering and $T$ comes with a cyclic ordering. Without loss of generality we may take $S=[n]:=\{1,2,\dots,n\}$ with its usual ordering and $T=[k]$ with the mod $k$ cyclic ordering. The action is generated by an operation $\mathbf{w}:\mathcal{F}\rightarrow\mathcal{F}$ (“the whirling map”) that is in turn defined as the composition of simpler maps $w_{i}:\mathcal{F}\rightarrow\mathcal{F}$ (“whirling at ii”, with $i$ in $[n]$) that repeatedly add $1\bmod k$ to the value of $f$ at $i$ until we get a function in $\mathcal{F}$. Our main results are examples of the homomesy [^1] Greek for “same middle” phenomenon, introduced by the second and third authors [PR15] and defined as follows.
@@ 1. Introduction
The homomesy phenomenon is quite widespread throughout combinatorial dynamical systems. See [Rob16] for a survey article. Since it is a new area of research over the past roughly 15 years, there is still much to learn about it, including the best techniques for proving homomesy, what types of actions and statistics commonly yield homomesy, and what consequences can arise from homomesy.
Early on, almost all proven homomesy results were for actions whose order was straightforwardly computable in general and relatively small compared to the size of the ground set $\mathcal{S}$. So it is rarer and notable to find homomesies for actions with unpredictable orbit sizes, as is the case with the maps in this paper, which also generally have unknown order. Other such examples can be found in [EFG+16, ELM+24]. In some cases, a statistic which is always an integer has a non-integer average across every orbit, which means all of the orbit sizes must be multiples of the constant average’s denominator. This consequence of homomesy appears in Corollary 2.21.
The whirling maps $w_{i}$ are a generalization of the toggle maps that are heavily studied in dynamical algebraic combinatorics. Toggles were first introduced as a way of writing the rowmotion map as a composition of simpler operations on order ideals [CF95]. Striker later generalized them to a much wider variety of settings [Str18]. There have been many interesting homomesy and periodicity phenomena discovered for actions defined in terms of toggles; see e.g., [PR15, Rob16, Had21, EFG+16, JR18, ELM+24].
In recent work, Plante and the third author consider a whirling action on the set of $k$-bounded $P$-partitions of a poset $P$. Following in the footsteps of Haddadan [Had21], they prove this to be in equivariant bijection with rowmotion acting on order ideals of the poset $P\times[k]$ for any finite poset $P$. They then leverage this to obtain periodicity and homomesy results for rowmotion acting on the poset ${\sf V}\times[k]$, which has surprisingly good dynamical properties [PR24]. A key point here is that beyond the intrinsic interest of whirling as a generalization of toggling, its study can facilitate our understanding of rowmotion questions that arise independently. A related construction is discussed in [DTW03]; we have borrowed the term “whirling” from the paper’s notion of “whirling tours”.
@@
1.1. Acknowledgements
We have benefitted from useful discussions with Yue Cai, David Einstein, Max Glick, Shahrzad Haddadan, Sam Hopkins, Thomas McConville, Matthew Plante, Elizabeth Sheridan Rossi, Richard Stanley, Jessica Striker, and Nathan Williams. Computations leading to the initial conjectures were done in Sage [S+16]. We also thank two anonymous reviewers who have provided helpful feedback.
2. Whirling on m-injections and m-surjections between finite sets
@@
2.1. The whirling map on m-injections and m-surjections
@@
We denote the set of all $m$-injective functions from $[n]$ to $[k]$ by $\operatorname{Inj}_{m}(n,k)$ and the set of all $m$-surjective functions from $[n]$ to $[k]$ by $\operatorname{Sur}_{m}(n,k)$. We write a function $f:[n]\rightarrow[k]$ in one-line notation as $f(1)f(2)\cdots f(n)$. For example, the function $f(1)=5$, $f(2)=1$, $f(3)=2$, $f(4)=1$ is written $f=5121$. (All of our examples will have codomain $[k]$ with $k\leq 9$.) Note that $m$-injections are not injective in general. They are called “width-$m$ restricted functions” in [Wal], which discusses their enumeration.
@@
For a family of functions $\mathcal{F}$ from $[n]$ to $[k]$, we define the (left-to-right) whirling map, denoted $\mathbf{w}:\mathcal{F}\rightarrow\mathcal{F}$, to be the map that whirls at indices $1,2,\dots,n$ in that order. So $\mathbf{w}=w_{n}\circ\cdots\circ w_{2}\circ w_{1}$.
@@
2.2. Orbit board terminology and notation
In this subsection, we introduce terminology and notation that applies to whirling on any family of functions $\mathcal{F}$ from $[n]$ to $[k]$ that we will use on various homomesy proofs. We use a generalization of the tuple boards of Haddadan; see [Had21].
@@ 2.2. Orbit board terminology and notation
Figure 1. The orbit under the action of $\mathbf{w}$ on $\operatorname{Inj}_{1}(3,6)$ containing $f=621$.
@@
2.3. The proof of Theorem 2.10 for injections
@@
Proof of Theorem 2.10 for $\mathcal{F}=\operatorname{Inj}_{1}(n,k)$.
@@
2.4. The proof of Theorem 2.10 for m-injections
@@
Figure 2. Two different ways to partition the same $\mathbf{w}$-orbit of $\operatorname{Inj}_{2}(4,4)$ into $[4]$-chunks.
@@
Proof of Theorem 2.10 for $\mathcal{F}=\operatorname{Inj}_{m}(n,k)$.
@@
Figure 3. In the left diagram, we cannot complete the purple chunk because the only position containing 3 within the next $n=4$ positions after the purple chunk’s 2 is in the brown chunk already. Thus, we have to take the end of the brown chunk starting from the 3, place it in the purple chunk (middle), and then complete the brown chunk (right).
@@
2.5. The proof of Theorem 2.10 for surjections
@@
If (P,[1,n1])(P,[1,n-1]) does not contain $i$ in any position, then $(w_{j-1}\circ\cdots\circ w_{1})(f)$ only contains $i$ as an output once. So to maintain surjectivity, $((w_{j}\circ\cdots\circ w_{1})(f))(j)=i$, and hence $(P,n)$ contains the value $i$.
@@ 1 $\normalsize{{\color{black}4}}$⃝ 3 $\normalsize{{\color{black}3}}$⃝ 1 2
Figure 4. An example $\mathbf{w}$-orbit of $\operatorname{Sur}_{1}(8,4)$ containing $f=31114424$. For each stoplight $P$ containing $i$, the last position in (P,[1,n1])(P,[1,n-1]) that contains $i+1$ (which is a stoplight by Lemma 2.20) is shown in the same color.
@@
Proof of Theorem 2.10 for $\mathcal{F}=\operatorname{Sur}_{1}(n,k)$.
@@ 2.6. Consequences of the homomesy
Corollary 2.21 gives an example where we can use the homomesy to prove a result that neither mentions homomesy nor the statistic that is homomesic. As the sizes of the orbits are generally unpredictable, this is the only known way to prove this. This is one of several instances where homomesy has been used to determine divisibility properties of a map; see  [EFG+16, Corollary 4.8] for another.
@@
Functions from a set $E$ to $[2]$ are in a natural correspondence with subsets of $E$ via indicator functions. In this case, whirling at index $i$ is equivalent to toggling at index $i$. Toggle groups were first introduced by Cameron and Fon-der-Flaass in the specific setting of order ideals of a poset [CF95] and generalized more recently by Striker [Str18] via the definition below.
Definition 2.22 ([Str18]).
@@
Corollary 2.23 is one of many homomesic statistics in toggle group actions defined as a product of every toggle each used exactly once. Other such results include rowmotion and promotion on order ideals and antichains of various posets, as well as actions on independent sets of graphs and on noncrossing partitions (including the Kreweras complement) [PR15, Rob16, Had21, EFG+16, JR18, ELM+24].
2.7. Lifting whirling to the piecewise-linear realm
@@
We have tested several random orbits and we conjecture that the homomesy in Corollary 2.23 extends to $\mathcal{P}_{r}(n)$. There is one caveat though. A vector with irrational entries probably does not always yield a finite orbit. Thus, we refer to the notion of asympotic homomesy, discussed in [Vor18, Jos19, DL24], and which is clearly equivalent to the previous definition of homomesy when all of the orbits are finite.
Definition 2.27 ([Vor18, Definition 5.3.1]).
Suppose we have a set $\mathcal{S}$, a map $w:\mathcal{S}\rightarrow\mathcal{S}$, and a function (statistic) $f:\mathcal{S}\rightarrow\mathbb{R}$. If there exists $c\in\mathbb{R}$ such that for every $x\in\mathcal{S}$,
@@
It is not uncommon for homomesies in the combinatorial realm to lift to the piecewise-linear realm; see [EP21, Jos19, HJ22, Hop22]. However, as far as we know, the only homomesies shown to lift to the piecewise-linear realm are on the order polytope or chain polytope of a poset. So Conjecture 2.28 is interesting as it would be an example outside of a poset.
3. Whirling on restricted growth words
3.1. Basic definitions, properties, and main result
@@
Let $\left\{n\atop k\right\}$ denote the number of partitions of $[n]$ with exactly $k$ blocks and $B_{n}$ denote the total number of partitions of $[n]$. These are called the Stirling numbers of the second kind and Bell numbers respectively; e.g., see [Sta11, §1.9]. RG-words were first introduced by Hutchinson [Hut63] and have been studied more recently by Cai and Readdy [CR17] for their connections to the $q$ and “qq-(1+q)(1+q)” analogues for Stirling numbers of the second kind.
@@
Figure 5. The two $\mathbf{w}$-orbits for $\mathcal{F}=\operatorname{RG}(5,3)$. The left orbit has length 16 and the right orbit has length 9. Notice that in the left orbit, there are the same numbers (four each) of functions $h$ satisfying any of the conditions $h(2)=1$, $h(3)=1$, $h(4)=1$, or $h(5)=1$. In the right orbit, there are also the same numbers (two each) of functions $h$ satisfying any of the conditions $h(2)=1$, $h(3)=1$, $h(4)=1$, or $h(5)=1$. This illustrates Theorem 3.8 for the case $n=5$, $k=3$, and $\pi$ is the identity.
@@
3.2. Proof of homomesy for whirling restricted growth words
@@
(1) $g(i)<g(i+1)\text{ and }(\mathbf{w}_{\pi}(g))(i)\geq(\mathbf{w}_{\pi}(g))(i+1).$
@@
3.3. Whirling on noncrossing RG-words

### v6

@@ Abstract
keywords:
dynamical algebraic combinatorics, homomesy, injection, noncrossing partition, RG-word, set partition, surjection, toggles, toggle group, whirling
1 Introduction
In this paper, we explore the orbits associated with an action that can be defined on any family $\mathcal{F}$ of functions $f$ from one finite set $S$ to another finite set $T$, where $S$ comes with a linear ordering and $T$ comes with a cyclic ordering. Without loss of generality we may take $S=[n]:=\{1,2,\dots,n\}$ with its usual ordering and $T=[k]$ with the mod $k$ cyclic ordering. The action is generated by an operation $\mathbf{w}:\mathcal{F}\rightarrow\mathcal{F}$ (“the whirling map”) that is in turn defined as the composition of simpler maps $w_{i}:\mathcal{F}\rightarrow\mathcal{F}$ (“whirling at ii”, with $i$ in $[n]$) that repeatedly add $1\bmod k$ to the value of $f$ at $i$ until we get a function in $\mathcal{F}$. Our main results are examples of the homomesy [^1] Greek for “same middle” phenomenon, introduced by the second and third authors Propp and Roby (2015) and defined as follows.
@@ 1 Introduction
The homomesy phenomenon is quite widespread throughout combinatorial dynamical systems. See Roby (2016) for a survey article. Since it is a new area of research over the past roughly 15 years, there is still much to learn about it, including the best techniques for proving homomesy, what types of actions and statistics commonly yield homomesy, and what consequences can arise from homomesy.
Early on, almost all proven homomesy results were for actions whose order was straightforwardly computable in general and relatively small compared to the size of the ground set $\mathcal{S}$. So it is rarer and notable to find homomesies for actions with unpredictable orbit sizes, as is the case with the maps in this paper, which also generally have unknown order. Other such examples can be found in Einstein et al. (2016); Elder et al. (2024). In some cases, a statistic which is always an integer has a non-integer average across every orbit, which means all of the orbit sizes must be multiples of the constant average’s denominator. This consequence of homomesy appears in Corollary 2.21.
The whirling maps $w_{i}$ are a generalization of the toggle maps that are heavily studied in dynamical algebraic combinatorics. Toggles were first introduced by Cameron and Fon-Der-Flaass (1995) as a way of writing the rowmotion map as a composition of simpler operations on order ideals. Striker (2018) later generalized them to a much wider variety of settings. There have been many interesting homomesy and periodicity phenomena discovered for actions defined in terms of toggles; see e.g., Propp and Roby (2015); Roby (2016); Haddadan (2021); Einstein et al. (2016); Joseph and Roby (2018); Elder et al. (2024).
In recent work, Plante and Roby (2024) consider a whirling action on the set of $k$-bounded $P$-partitions of a poset $P$. Following in the footsteps of Haddadan (2021), they prove this to be in equivariant bijection with rowmotion acting on order ideals of the poset $P\times[k]$ for any finite poset $P$. They then leverage this to obtain periodicity and homomesy results for rowmotion acting on the poset ${\sf V}\times[k]$, which has surprisingly good dynamical properties. A key point here is that beyond the intrinsic interest of whirling as a generalization of toggling, its study can facilitate our understanding of rowmotion questions that arise independently. A related construction is discussed in Dumitriu et al. (2003); we have borrowed the term “whirling” from the paper’s notion of “whirling tours”.
@@
2 Whirling on m-injections and m-surjections between finite sets
@@
2.1 The whirling map on m-injections and m-surjections
@@
We denote the set of all $m$-injective functions from $[n]$ to $[k]$ by $\operatorname{Inj}_{m}(n,k)$ and the set of all $m$-surjective functions from $[n]$ to $[k]$ by $\operatorname{Sur}_{m}(n,k)$. We write a function $f:[n]\rightarrow[k]$ in one-line notation as $f(1)f(2)\cdots f(n)$. For example, the function $f(1)=5$, $f(2)=1$, $f(3)=2$, $f(4)=1$ is written $f=5121$. (All of our examples will have codomain $[k]$ with $k\leq 9$.) Note that $m$-injections are not injective in general. They are called “width-$m$ restricted functions” by Walsh , who discusses their enumeration.
@@
For a family of functions $\mathcal{F}$ from $[n]$ to $[k]$, we define the (left-to-right) whirling map $\mathbf{w}:\mathcal{F}\rightarrow\mathcal{F}$ to be the map that whirls at indices $1,2,\dots,n$ in that order. So $\mathbf{w}=w_{n}\circ\cdots\circ w_{2}\circ w_{1}$.
@@
2.2 Orbit board terminology and notation
In this subsection, we introduce terminology and notation that applies to whirling on any family of functions $\mathcal{F}$ from $[n]$ to $[k]$ that we will use on various homomesy proofs. We use a generalization of the tuple boards of Haddadan (2021).
@@ 2.2 Orbit board terminology and notation
Figure 1: The orbit under the action of $\mathbf{w}$ on $\operatorname{Inj}_{1}(3,6)$ containing $f=621$.
@@
2.3 The proof of Theorem 2.10 for injections
@@
of Theorem 2.10 for $\mathcal{F}=\operatorname{Inj}_{1}(n,k)$.
@@
2.4 The proof of Theorem 2.10 for m-injections
@@
Figure 2: Two different ways to partition the same $\mathbf{w}$-orbit of $\operatorname{Inj}_{2}(4,4)$ into $[4]$-chunks.
@@
of Theorem 2.10 for $\mathcal{F}=\operatorname{Inj}_{m}(n,k)$.
@@
Figure 3: In the left diagram, we cannot complete the purple chunk because the only position containing 3 within the next $n=4$ positions after the purple chunk’s 2 is in the brown chunk already. Thus, we have to take the end of the brown chunk starting from the 3, place it in the purple chunk (middle), and then complete the brown chunk (right).
@@
2.5 The proof of Theorem 2.10 for surjections
@@
If (P,[1,n1])(P,[1,n-1]) does not contain $i$ in any position, then $(w_{j-1}\circ\cdots\circ w_{1})(f)$ only contains $i$ as an output once. So to maintain surjectivity, $((w_{j}\circ\cdots\circ w_{1})(f))(j)=i$, and thus $(P,n)$ contains the value $i$.
@@ 1 $\normalsize{{\color{black}4}}$⃝ 3 $\normalsize{{\color{black}3}}$⃝ 1 2
Figure 4: An example $\mathbf{w}$-orbit of $\operatorname{Sur}_{1}(8,4)$ containing $f=31114424$. For each stoplight $P$ containing $i$, the last position in (P,[1,n1])(P,[1,n-1]) that contains $i+1$ (which is a stoplight by Lemma 2.20) is shown in the same color.
@@
of Theorem 2.10 for $\mathcal{F}=\operatorname{Sur}_{1}(n,k)$.
@@ 2.6 Consequences of the homomesy
Corollary 2.21 gives an example where we can use the homomesy to prove a result that neither mentions homomesy nor the statistic that is homomesic. As the sizes of the orbits are generally unpredictable, this is the only known way to prove this. This is one of several instances where homomesy has been used to determine divisibility properties of a map; see  (Einstein et al., 2016, Corollary 4.8) for another.
@@
Functions from a set $E$ to $[2]$ are in a natural correspondence with subsets of $E$ via indicator functions. In this case, whirling at index $i$ is equivalent to toggling at index $i$. Toggle groups were first introduced by Cameron and Fon-Der-Flaass (1995) in the specific setting of order ideals of a poset, and generalized more recently by Striker (2018) via the definition below.
Definition 2.22 (Striker (2018)).
@@
Corollary 2.23 is one of many homomesic statistics in toggle group actions defined as a product of every toggle each used exactly once. Other such results include rowmotion and promotion on order ideals and antichains of various posets, as well as actions on independent sets of graphs and on noncrossing partitions (including the Kreweras complement), see Propp and Roby (2015); Roby (2016); Haddadan (2021); Einstein et al. (2016); Joseph and Roby (2018); Elder et al. (2024).
2.7 Lifting whirling to the piecewise-linear realm
@@
We have tested several random orbits and we conjecture that the homomesy in Corollary 2.23 extends to $\mathcal{P}_{r}(n)$. There is one caveat though. A vector with irrational entries probably does not always yield a finite orbit. Thus, we refer to the notion of asympotic homomesy, discussed in Vorland (2018); Joseph (2019); Defant and Lin (2024), and which is clearly equivalent to the previous definition of homomesy when all of the orbits are finite.
Definition 2.27.
(Vorland, 2018, Definition 5.3.1). Suppose we have a set $\mathcal{S}$, a map $w:\mathcal{S}\rightarrow\mathcal{S}$, and a function (statistic) $f:\mathcal{S}\rightarrow\mathbb{R}$. If there exists $c\in\mathbb{R}$ such that for every $x\in\mathcal{S}$,
@@
It is not uncommon for homomesies in the combinatorial realm to lift to the piecewise-linear realm; see Einstein and Propp (2021); Joseph (2019); Hopkins and Joseph (2022); Hopkins (2022). However, as far as we know, the only homomesies shown to lift to the piecewise-linear realm are on the order polytope or chain polytope of a poset. So Conjecture 2.28 is interesting as it would be an example outside of a poset.
3 Whirling on restricted growth words
3.1 Basic definitions, properties, and main result
@@
Let $\left\{n\atop k\right\}$ denote the number of partitions of $[n]$ with exactly $k$ blocks and $B_{n}$ denote the total number of partitions of $[n]$. These are called the Stirling numbers of the second kind and Bell numbers respectively; e.g., see (Stanley, 2011, §1.9). RG-words were first introduced by Hutchinson (1963) and have been studied more recently by Cai and Readdy (2017) for their connections to the $q$ and “qq-(1+q)(1+q)” analogues for Stirling numbers of the second kind.
@@
Figure 5: The two $\mathbf{w}$-orbits for $\mathcal{F}=\operatorname{RG}(5,3)$. The left orbit has length 16 and the right orbit has length 9. Notice that in the left orbit, there are the same numbers (four each) of functions $h$ satisfying any of the conditions $h(2)=1$, $h(3)=1$, $h(4)=1$, or $h(5)=1$. In the right orbit, there are also the same numbers (two each) of functions $h$ satisfying any of the conditions $h(2)=1$, $h(3)=1$, $h(4)=1$, or $h(5)=1$. This illustrates Theorem 3.8 for the case $n=5$, $k=3$, and $\pi$ is the identity.
@@
3.2 Proof of homomesy for whirling restricted growth words
@@
$g(i)<g(i+1)\text{ and }(\mathbf{w}_{\pi}(g))(i)\geq(\mathbf{w}_{\pi}(g))(i+1).$ (1)
@@
3.3 Whirling on noncrossing RG-words
@@
Acknowledgements.
We have benefitted from useful discussions with Yue Cai, David Einstein, Max Glick, Shahrzad Haddadan, Sam Hopkins, Thomas McConville, Matthew Plante, Elizabeth Sheridan Rossi, Richard Stanley, Jessica Striker, and Nathan Williams. Computations leading to the initial conjectures were done in Sage (Stein et al. (2016)). We also thank two anonymous reviewers who have provided helpful feedback.